OpenStudy (anonymous):

Chris entered a hot dog eating contest. The table shows how many hot dogs remained in his tray at different times. How many hot dogs did Chris eat per minute? Enter the answer as a decimal. [_____] Chris’s Contest Minutes 2 4 6 8 Hot dogs remaining 49 38 27 16

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

plz help sorry i am doing two subjects

OpenStudy (alexandervonhumboldt2):

49-38=11 11/2=5.5 i guees

OpenStudy (anonymous):

well you never failed me Alexander

OpenStudy (alexandervonhumboldt2):

he eats 11 dogs per 2 minutes so 1 minute 11/2=5.5

OpenStudy (anonymous):

oh! i get it

OpenStudy (anonymous):

sort of not really good at this stuff

OpenStudy (alexandervonhumboldt2):

ok